OpenStudy (anonymous):

help pls how to differentiate (2x+1)/2sqrt(x^2+x-2)

OpenStudy (mathmale):

I'm going to assume that you mean the following:\[\frac{ 2x+1 }{ 2\sqrt(x^2+x-2) }\]

OpenStudy (mathmale):

Since this is a quotient, apply the quotient rule for differentiation.

OpenStudy (anonymous):

yep thats the problem

OpenStudy (mathmale):

The formula for the derivative of the quotient y=u/v is\[\frac{ dy }{ dx }=\frac{ v \frac{ du }{ dx }-u \frac{ dv }{ dx } }{ v^2 }\]

OpenStudy (anonymous):

Another approach you could consider, though not necessarily as simple as just using the quotient rule. Notice that the numerator is the derivative of the highlighted term in the denominator: \[\frac{ 2x+1 }{ 2\sqrt{\color{red}{x^2+x-2}}}\] So if you were to let \(t=x^2+x-2\), you'd have \[\frac{t'}{ 2\sqrt t}=\frac{1}{2}t^{-1/2}t'\]

OpenStudy (mathmale):

If u = 2x+1, what is du/dx ? If v = 2 Sqrt(x^2+2-2), what is dv/dx ?

OpenStudy (mathmale):

the approach suggested by sithsandgiggles is an excellent one for this particular problem.

OpenStudy (anonymous):

u = 2 and v=(x^2+x-2)(2x+1) right?

OpenStudy (mathmale):

Actually, u=2x+1, so that du/dx = 2. (Label every quantity properly.) v=2Sqrt(x^2+x-2). What is dv / dx ?

OpenStudy (mathmale):

Note that the quantity x^2 + x - 2 has the derivative 2x + 1. Thus, you could re-write \[\frac{ 2x+1 }{ 2\sqrt{\color{red}{x^2+x-2}}}\]